# Efficient Fairness-Performance Pareto Front Computation

### v1

@@ Abstract
There is a well known intrinsic trade-off between the fairness of a representation and the performance of classifiers derived from the representation. Due to the complexity of optimisation algorithms in most modern representation learning approaches, for a given method it may be non-trivial to decide whether the obtained fairness-performance curve of the method is optimal, i.e., whether it is close to the true Pareto front for these quantities for the underlying data distribution.
In this paper we propose a new method to compute the optimal Pareto front, which does not require the training of complex representation models. We show that optimal fair representations possess several useful structural properties, and that these properties enable a reduction of the computation of the Pareto Front to a compact discrete problem. We then also show that these compact approximating problems can be efficiently solved via off-the shelf concave-convex programming methods.
Since our approach is independent of the specific model of representations, it may be used as the benchmark to which representation learning algorithms may be compared. We experimentally evaluate the approach on a number of real world benchmark datasets.
@@ 1 Introduction
Fair representations are a central topic in the field of Fair Machine Learning, Mehrabi et al., (2021), Pessach and Shmueli, (2022),Chouldechova and Roth, (2018). Since their introduction in Zemel et al., (2013), Fair representations have been extensively studied, giving rise to a variety of approaches based on a wide range of modern machine learning methods, such GANs, variational auto encoders, numerous variants of Optimal Transport methods, and direct variational formulations. See the papers Feldman et al., (2015), Edwards and Storkey, (2015),Louizos et al., (2015),Madras et al., (2018), Gordaliza et al., (2019); Zehlike et al., (2020), Song et al., (2019), Du et al., (2020), for a sample of existing methods.
@@
For a given a representation learning problem and a target classification problem, since the fairness constraints reduce the space of feasible classifiers, the best possible classification performance will usually be lower as the fairness constraint becomes stronger. This phenomenon is known as the Fairness-Performance trade-off. Assume that we have fixed a way to measure fairness. Then for a given representation learning method, one is often interested in the fairness-performance curve $(\gamma,E(\gamma))$. Here, $\gamma$ is the fairness level, and $E(\gamma)$ is the classification performance of the method at that level. The curve $(\gamma,E(\gamma))$ where $E(\gamma)$ is the best possible performance over all representations and classifiers under the constraint is known as the Fairness-Performance Pareto Front.
@@
As indicated by the above discussion, representation learning methods typically involve models with high dimensional parameter spaces, and complex, possibly constrained non-convex optimisation algorithms. As such, these methods may be prone to local minima, and may be sensitive to a variety of hyper-parameters, such as architecture details, learning rates, and even initializations. While the representations produced by such methods may often be useful, it nevertheless may be difficult to decide whether their associated Fairness-Performance curve is close to the true Pareto Front.
@@
In this paper we propose a new method to compute the optimal Pareto front, which does not require the training of complex fair representation models. In other words, we show that, perhaps somewhat surprisingly, the computation of the Pareto Front can be decoupled from that of the representation, and only relies on learning of much simpler, unconstrained classifiers on the data. To achieve this, we first show that the optimal fair representations satisfy a number of structural properties. While these properties may be of independent interest, here we use them to express the points on the Pareto Front as the solutions of small discrete optimisation problems. The problems fall into the category of concave minimisation problems Benson, (1995), which have been extensively studied, and can be efficiently solved using modern dedicated optimisation frameworks, Shen et al., (2016).
@@
We now describe the results in more detail. Let $X$ and $A$ denote the data features and the sensitive attribute, respectively. The sensitive attribute is a part of the data, and we assume that $A$ is binary. The features $X$ will be taking values in $\mathbb{R}^{d}$. In addition, we have a target variable $Y$, taking values in a finite set $\mathcal{Y}$. That is, we are interested in representations that maximise the performance of prediction of $Y$, under the fairness constraint. The representation is denoted by $Z$, and is typically expressed by constructing the conditional distributions $\mathbb{P}\left(Z|X=x,A=a\right)$, for $a\in\left\{0,1\right\}$. The problem setting is illustrated in Figure 1.
Figure 1: Fair Representation Problem Setting
In Section 4 we introduce the invertibility theorem. This result states, informally, that in an optimal representation, every point $z$ can be generated from at most two $x$’s, one given $A=0$ and another given $A=1$. That is, there are unique $u,v$ such that $\mathbb{P}\left(Z=z|X=u,A=0\right)\neq 0$ and $\mathbb{P}\left(Z=z|X=v,A=1\right)\neq 0$. This in turn hints at a possibility that in an optimal representation, the space of $z$ may be indexed by the pairs $(u,v)$, where $u,v$ denote $x$’s corresponding to $A=0$ and $A=1$ as above. Indeed, an additional result, the Uniform Approximation Lemma D.1, deferred to Supplementary D due to space constraints, implies that for every pair $(u,v)$, one needs to have only a small number of $z$’s associated with it, in order to approximate all possible representations. Together, these results bound the number of points needed to model all optimal representations, if $X$ has finitely many values.
Next, in Section 5.1 we show that instead of discretising $X$ as suggested above, one may in fact discretise a much smaller space, i.e., $\Delta_{\mathcal{Y}}$—the space of all possible distributions of $Y$. As an example, if $Y$ is binary, this space is the interval $[0,1]$. This reduction is possible due to the factorisation property of representations. That is, we show that for the purposes of the computation of the Pareto front, any representation may be represented as a composition (thus the factorisation terminology) of a mapping into $\Delta_{\mathcal{Y}}$, and a representation of $\Delta_{\mathcal{Y}}$.
@@
Based on these results, in Section 5.2 we introduce MIFPO, a discrete optimisation problem for the computation of the Pareto front, and in Section 5.3 we provide details on estimating the problem parameters from the data. The MIFPO problem is a concave minimisation problem with linear constraints, and we solve it using the disciplined convex-concave programming framework, DCCP, Shen et al., (2016). It is worth noting that for the special case of perfect fairness, i.e., the point $\gamma=0$ on the curve, the MIFPO problem is in fact a linear programming problem, solvable by standard means, and closely related to the Optimal Transportation problem.
We illustrate MIFPO by evaluating it on a number of standard real world fairness benchmark datasets. We also compare its fairness-performance curve to that of FNF, Balunović et al., 2022a , a recent representation learning approach which yields representations with guaranteed fairness constraint satisfaction. As expected, MIFPO produces tighter fairness-performance curves, indicating that FNF still has space for improvement. This suggests that MIFPO may also be used as an analysis method in the construction of full representation learning algorithms. In a similar fashion, in Supplementary Material Section H, we show that MIFPO may also be used to evaluate the fairness-performance Pareto front of fair classifiers, where fairness is measured by statistical parity. We evaluate two widely used fair classification methods and find that their Pareto front is significantly below that of MIFPO.
@@
The rest of this paper is organised as follows: Section 2 discusses the literature. Section 3 introduces the formal problem setting. The Inveribility Theorem is discussed in Section 4, and the details of the MIFPO construction are given in Section 5. Experimental results are presented in Section 6, and we conclude the paper in Section 7. All proofs are provided in the Supplementary Material.
@@ 2 Literature and Prior Work
As mentioned in the previous Section, there exists a variety of approaches to the construction of fair representations. We refer to the book Barocas et al., (2023), and surveys Mehrabi et al., (2021),Du et al., (2020), for an overview. Notably, the work Song et al., (2019) introduces a variational framework that unifies many of the existing methods under a single optimisation procedure. This work also explicitly deals with the difficulty of obtaining useful fairness-performance trade-offs, although such considerations are present in practically all fairness related work.
In a recent paper, Balunović et al., 2022a , it was observed that if one knows the density of the source data, $\mathbb{P}\left(X|A=a\right)$, $a\in\left\{0,1\right\}$, and uses models with computable change of variables, such as the normalising flows, Rezende and Mohamed, (2015), then one can verify the total variation fairness criterion (see (3)), with high probability. This enables one to provide guarantees on the fairness of a representation, which is clearly an important feature in applications. Such guarantees also allow result comparison between different methods. Indeed, if a given constraint is not guaranteed in one of the methods, then the performance comparison may be unfair. Due to this reason, in the experiments Section 6 we compare our own results to those of Balunović et al., 2022a . It may be worthwhile to note, however, that the assumption of knowing the densities $\mathbb{P}\left(X|A=a\right)$, made in Balunović et al., 2022a , is quite significant. This is due to the fact that such information is normally not given, and the estimation of densities from data is generally a difficult problem. We remark that our approach does not rely on this assumption.
3 Problem Setting
@@
Let $A$ be a binary sensitive variable, and let $X$ be an additional feature random variable, with values in $\mathbb{R}^{d}$. Assume also that there is a target variable $Y$ with finitely many values, jointly distributed with $X,A$.
@@
A representation $Z$ of $(X,A)$ is defined as a random variable taking values in some space $\mathbb{R}^{d^{\prime}}$, with (i) distribution given through $\mathbb{P}_{\theta}\left(Z|X,A\right)$, where $\theta$ are the parameters of the representation, and (ii) such that $Z$ is independent of the rest of the variables of the problem conditioned on $(X,A)$. In particular, we have
@@
[…]
For $a\in\left\{0,1\right\}$, let $\mu_{a}$ be the distribution of $Z$ given $A=a$, i.e. $\mu_{a}(\cdot):=\mathbb{P}\left(Z=\cdot|A=a\right)$. For $\gamma\geq 0$, we say that the representation $Z$ is $\gamma$-fair iff
[…]
Next, we describe the measurement of information loss in $Y$ due to the representation. Let $h:\Delta_{\mathcal{Y}}\rightarrow\mathbb{R}$ be a continuous and concave function on the set of probability distributions on $\mathcal{Y}$, $\Delta_{\mathcal{Y}}$. The quantity $h(\mathbb{P}\left(Y|X=x\right))$ will measure the best possible prediction accuracy of $Y$ conditioned on $X=x$, for varying $x$. As an example, consider the case of binary $Y$, $\mathcal{Y}=\left\{0,1\right\}$. Every point in $\Delta_{\mathcal{Y}}$ can be written as $(p,1-p)$ for $p\in[0,1]$, and we may choose $h$ to be the optimal binary classification error,
$h((1-p,p))=min(p,1-p).$ (4)
@@
In supplementary material Section A we show that for any representation, $\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\geq\mathbb{E}_{x\sim X}h(\mathbb{P}\left(Y|X=x\right))$, i.e. representations generally decrease the performance. Technically, this happens due to the concavity of $h$, and this result is an extension of the classical information processing inequality, Cover and Thomas, (2012). To intuitively see why representations decrease performance, note that as illustrated in Figure 1, the distribution $\mathbb{P}\left(Y|Z=z\right)$ is a mixture of various distributions $\mathbb{P}\left(Y|X=x,A=a\right)$. Thus, for instance, even if all $\mathbb{P}\left(Y|X=x,A=a\right)$ are deterministic, but not all of the same value, $\mathbb{P}\left(Y|Z=z\right)$ will not be deterministic.
4 The Invertibility Theorem

### v2

@@ Abstract
There is a well known intrinsic trade-off between the fairness of a representation and the performance of classifiers derived from the representation. In this paper we propose a new method to compute the optimal Pareto front of this trade off. In contrast to the existing methods, this approach does not require the training of complex fair representation models.
Our approach is derived through three main steps: We analyze fair representations theoretically, and derive several structural properties of optimal representations. We then show that these properties enable a reduction of the computation of the Pareto Front to a compact discrete problem. Finally, we show that these compact approximating problems can be efficiently solved via off-the shelf concave-convex programming methods.
In addition to representations, we show that the new methods may also be used to directly compute the Pareto front of fair classification problems. Moreover, the proposed methods may be used with any concave performance measure. This is in contrast to the existing reduction approaches, developed recently in fair classification, which rely explicitly on the structure of the non-differentiable accuracy measure, and are thus unlikely to be extendable.
The approach was evaluated on several real world benchmark datasets and compares favorably to a number of recent state of the art fair representation and classification methods.
@@ 1 Introduction
Fair representations are a central topic in the field of Fair Machine Learning, Mehrabi et al., (2021), Pessach and Shmueli, (2022),Chouldechova and Roth, (2018). Since their introduction in Zemel et al., (2013), Fair representations have been extensively studied, giving rise to a variety of approaches based on a wide range of modern machine learning methods, such GANs, variational auto encoders, numerous variants of Optimal Transport methods, and direct variational formulations. See the papers Feldman et al., (2015), Madras et al., (2018), Gordaliza et al., (2019); Zehlike et al., (2020), Song et al., (2019), Du et al., (2020), Zhao and Gordon, (2022), Jovanović et al., (2023), Dehdashtian et al., (2024), for a sample of existing methods.
@@
For a given representation learning problem and a target classification problem, since the fairness constraints reduce the space of feasible classifiers, the best possible classification performance will usually be lower as the fairness constraint becomes stronger. This phenomenon is known as the Fairness-Performance trade-off. Assume that we have fixed a way to measure fairness. Then for a given representation learning method, one is often interested in the fairness-performance curve $(\gamma,E(\gamma))$. Here, $\gamma$ is the fairness level, and $E(\gamma)$ is the classification performance of the method at that level. The curve $(\gamma,E(\gamma))$ where $E(\gamma)$ is the best possible performance over all representations and classifiers under the constraint is known as the Fairness-Performance Pareto Front.
@@
As indicated by the above discussion, representation learning methods typically involve models with high dimensional parameter spaces, and complex, possibly constrained non-convex optimisation algorithms. As such, these methods may be prone to local minima and sensitivity to a variety of hyper-parameters, such as architecture details, learning rates, and even initializations. While the representations produced by such methods may often be useful, it nevertheless may be difficult to decide whether their associated Fairness-Performance curve is close to the true Pareto Front.
@@
In this paper we propose a new method to compute the optimal Pareto front, which does not require the training of complex fair representation models. In other words, we show that, perhaps somewhat surprisingly, the computation of the Pareto Front can be decoupled from that of the representation, and only relies on learning of much simpler, unconstrained classifiers on the data. To achieve this, we first show that the optimal fair representations satisfy a number of structural properties. While these properties may be of independent interest, here we use them to express the points on the Pareto Front as the solutions of small discrete optimisation problems. These problems, known as concave minimisation problems Benson, (1995), have been extensively studied and can be efficiently solved using modern dedicated optimisation frameworks, Shen et al., (2016).
@@
We now describe the results in more detail. Let $X\in\mathcal{X}$ and $A\in\mathcal{A}$ denote the data features and the sensitive attribute, respectively. We assume that $A$ is binary, while $X$ may take values in an arbitrary space $\mathcal{X}$, typically with $\mathcal{X}=\mathbb{R}^{d}$. In addition, we have a target variable $Y$, taking values in a finite set $\mathcal{Y}$. We are then interested in representations that maximise the performance of prediction of $Y$, under the fairness constraint. The representation is denoted by $Z$, and is typically expressed by constructing the conditional distributions $\mathbb{P}\left(Z|X=x,A=a\right)$, for $a\in\left\{0,1\right\}$. The problem setting is illustrated in Figure 1.
Figure 1: Fair Representation Problem Setting.
Our approach consists of three main steps: We first observe that one can map the data features $(x,a)\in\mathcal{X}\times\mathcal{A}$ to a much smaller space $\Delta_{\mathcal{Y}}$ of distributions on the set of label values $\mathcal{Y}$, without loosing any information necessary for the computation of the Pareto front. The mapping is done by means of the optimal Bayes classifier. This result is referred to as Factorization Lemma, Section 3.2, where the mapping is done via the optimal Bayes classifier. Similar arguments were recently implicitly used in the study of fair classification tradeoffs, (Xian et al.,, 2023; Wang et al.,, 2023), but were restricted to classification and to accuracy loss (Section 2).
The advantage of working with a small space such as $\Delta_{\mathcal{Y}}$ is that it can be easily discretized. For instance, if Y=0,1Y={0,1}, then we essentially have $\Delta_{\mathcal{Y}}=[0,1]$, which is descretised trivially. We note that other, data dependent discretisation schemes, such as clustering, maybe possible for problems involving highly multi label targets. Alternatively, one can also consider the dataset itself as a grid, a view that is typically taken by transportation based approaches, e.x. Gordaliza et al., (2019), Xian et al., (2023).
Next, assuming the data is discretized and finite, we ask how large the represntation space $\mathcal{Z}$ should be, in order to support both optimal performance and fairness? For instance, we believe the answer to the following question is not apriori obvious: Can representations on infinite spaces be approximated, in terms of performance and fairness, by representations on finite and bounded spaces $\mathcal{Z}$? These questions are addressed by the Invertibility Theorem, Section 3.3, which asserts that all optimal representations may be taken in a certain canonical form, which we term invertible. This result, in conjunction with an additional approximation lemma, is used in Section 4.1 to construct representations with any desired degree of approximation.
@@
Finally, based on these results in Section 4.1 we also introduce the MIFPO (Model Independent Fairness-Performance Optimization), a discrete optimisation problem that is essentially equivalent to a computation of the fairness-performance tradeoff on a discrete set. We show that in this situation MIFPO is a concave minimisation problem with linear constraints, and we solve it using the disciplined convex-concave programming framework, DCCP, Shen et al., (2016).
We evaluate our approach on standard fairness benchmark datasets and compare its fairness-performance curve to multiple state-of-the-art fair representation methods. We also compare MIFPO to the fairness-performance Pareto front of fair classifiers [^1] See Sections 2 and 3.2 for the relation between our representation framework and fair classification results.. As expected, MIFPO effectively serves as an upper bound on almost all other algorithms in both cases.
@@
The rest of this paper is organised as follows: Section 2 discusses the literature and related work. In Section 3 we discuss the theoretical results, including factorization and the Invertibility Theorem. The MIFPO problem construction and the full Pareto Front computation algorithm are provided in Section 4. Experimental results are presented in Section 5, and we conclude the paper in Section 6. All proofs are provided in the Supplementary Material.
@@ 2 Literature and Prior Work
We refer to the book Barocas et al., (2023), and surveys Mehrabi et al., (2021),Du et al., (2020), for a general overview of representations. Tradeoffs in particular where explicitly studied in Song et al., (2019), Balunović et al., 2022a , Zhao and Gordon, (2022), Jovanović et al., (2023),Dehdashtian et al., (2024), among others.
In this paper we use the total variation based fairness constraints, similarly to the line of work in Madras et al., (2018), Zhao and Gordon, (2022), Balunović et al., 2022a , Jovanović et al., (2023). Other constraints used in the literature include entropy based constraints, Song et al., (2019), or RKHS based independence test constraints, Dehdashtian et al., (2024).
As discussed earlier, the vast majority of the work above concentrates on finding neural network based fair representations via involved optimization schemes with possible local minima, which may be hard to analyze. This highlights the usefulness of our approach direct approach to the computation of the Pareto front, which has clear theoretical grounding, and in which sources of approximation error are well understood and may be controlled.
Relations between fairness and performance were studied in Zhao and Gordon, (2022). In particular, for perfectly fair representations, they derived lower bounds on the accuracy in terms of the difference of the base rates between the groups. However, this work did not introduce new algorithms for the computation of fair representations or of the associated Pareto front. The extension of the considerations in this paper to the full front was carried in Xian et al., (2023) for classification (see below).
The accuracy fairness tradeoff has also been extensively studied in the context of fair classification (without representations), see for instance Agarwal et al., (2018), Kim et al., (2020) , Alghamdi et al., (2022), Xian et al., (2023), Wang et al., (2023), for a sample of recent approaches. In particular, the papers Xian et al., (2023), Wang et al., (2023) are state of the art, and are also the most closely related to our methods, among the existing work.
Similarly to our approach, the analysis in these two papers starts with the estimation of the probabilities $\mathbb{P}\left(Y|X,A\right)$, which are then used to compute the constrained performance. However, the subsequent steps are different. Crucially, the analysis in both Xian et al., (2023) and Wang et al., (2023) relies critically on the properties of the accuracy as the performance metric. Consequently, it can not be extended to general concave performance measures, such as the standard (minus) log loss, for instance. Roughly speaking, in the appropriate sense, accuracy largely ignores classification probabilities. This allows the simple description of classifiers as small confusion matrices in Wang et al., (2023) (extending the approach of Kim et al., (2020)), and the restriction of the distributions to the vertices of the simplex in Xian et al., (2023). The special structure of accuracy is highlighted also in our Lemma 3.1, where we show that classifiers with accuracy may be effectively described by representations using only 2 points. We conclude that even when restricted to classification, our approach analyses a fundamentally more complex situation compared to previous work. On the other hand, Xian et al., (2023) and Wang et al., (2023) support non binary sensitive attributes and group labels, while such an extension for our methods is out of scope for this paper. A comparison of computational complexities for these algorithms may be found in Supplementary J.
3 Structure of Fair Representations
In this Section we describe several theoretical properties of fair representations. In Section 3.1 we introduce the problem setup and the necessary notation. In Section 3.2 we discuss relations to classification with accuracy loss and the factorization result, which allows to reduce the size of the representation space. The Invertibility Theorem is introduced in Section 3.3.
3.1 Problem Setting
@@
Let $A$ be a binary sensitive variable, and let $X$ be an additional feature random variable, with values in a set $\mathcal{X}$, typically with $\mathcal{X}=\mathbb{R}^{d}$. Assume also that there is a target variable $Y$ with finitely many values in a set $\mathcal{Y}$, jointly distributed with $X,A$.
@@
A representation $Z$ of $(X,A)$ is defined as a random variable taking values in some space $\mathcal{Z}$, with (i) distribution given through $\mathbb{P}_{\theta}\left(Z|X,A\right)$, where $\theta$ are the parameters of the representation, and (ii) such that conditioned on $(X,A)$, $Z$ is independent of the rest of the variables of the problem. In particular, we have
@@
$\left\|\mu-\nu\right\|_{TV}$ $=\frac{1}{2}\sup_{g\text{ s.t. }\left\|g\right\|_{\infty}\leq 1}\int g(x)\cdot\left[f_{\mu}(x)-f_{\nu}(x)\right]dx$ $=\frac{1}{2}\int\left|f_{\mu}(x)-f_{\nu}(x)\right|dx.$ (2)
Note that $\int\left|f_{\mu}(x)-f_{\nu}(x)\right|dx$ is in fact the $L_{1}$ distance, and the equivalence $\left\|\cdot\right\|_{TV}=\frac{1}{2}\left\|\cdot\right\|_{L_{1}}$ is well known, see Cover and Thomas, (2012).
For $a\in\left\{0,1\right\}$, let $\mu_{a}$ be the distribution of $Z$ given $A=a$, i.e. $\mu_{a}(\cdot):=\mathbb{P}\left(Z=\cdot|A=a\right)$. We denote the distance induced by the representation as $D_{TV}(Z)=\left\|\mu_{0}-\mu_{1}\right\|_{TV}$, and for $\gamma\geq 0$, we say that the representation $Z$ is $\gamma$-fair iff
$D_{TV}(Z)=\left\|\mu_{0}-\mu_{1}\right\|_{TV}\leq\gamma\hskip 5.69054pt\hskip 5.69054pt\hskip 5.69054pt\hskip 5.69054pt\text{(Fairness Condition)}.$ (3)
@@
Next, we describe the measurement of information loss in $Y$ due to the representation. Let $h:\Delta_{\mathcal{Y}}\rightarrow\mathbb{R}$ be a continuous and concave function on the set of probability distributions on $\mathcal{Y}$, $\Delta_{\mathcal{Y}}$. The quantity $h(\mathbb{P}\left(Y|X=x\right))$ will measure the best possible prediction accuracy of $Y$ conditioned on $X=x$, for varying $x$. As an example, consider the case of binary $Y$, $\mathcal{Y}=\left\{0,1\right\}$. Every point in $\Delta_{\mathcal{Y}}$ can be written as (p,1p)(p,1-p) for $p\in[0,1]$, and we may choose $h$ to be the optimal binary classification error,
h((1p,p))=min(p,1p).h((1-p,p))=min(p,1-p). (4)
@@
In supplementary material Section A we show that for any representation, $\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\geq\mathbb{E}_{x\sim X}h(\mathbb{P}\left(Y|X=x\right))$, i.e. representations generally decrease or maintain the performance.
3.2 Classification and Factorization
In this section we show that the Pareto front of binary classifiers, with accuracy performance and statistical parity fairness measure, can be computed from the Pareto front of representations with total variation fairness measure. In fact, Lemma 3.1 below states that both Pareto fronts amount to the same curve. As discussed in Section 1, this equivalence implies that MIFPO can be used to evaluate fair classifiers, in addition to fair representations.
For a binary classifier $\hat{Y}$ of $Y$, with $(X,A)$ as features. The prediction error is defined as usual by $\epsilon(\hat{Y}):=\mathbb{P}\left(\hat{Y}\neq Y\right)$. The statistical parity of $\hat{Y}$ is defined as
$D_{SP}(\hat{Y}):=\left|\mathbb{P}\left(\hat{Y}=1|A=1\right)-\mathbb{P}\left(\hat{Y}=1|A=0\right)\right|.$ (6)
Lemma 3.1.
Let $\hat{Y}$ be a classifier of $Y$, let the representation uncertainty measure be given by (4). Then there is a representation given by a random variable $Z$ on a set $\mathcal{Z}$ with $\left|\mathcal{Z}\right|=2$, such that
[…]
$\epsilon(\hat{Y})\leq\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\text{ and }D_{SP}(\hat{Y})\leq\left\|\mu_{0}-\mu_{1}\right\|_{TV}.$ (8)
The Proof of Lemma 3.1 is presented in Supplementary Material Section I.
We now describe the Factorization result. Let $f^{*}:\mathcal{X}\times\mathcal{A}\rightarrow\Delta_{\mathcal{Y}}$ be the Bayes optimal classifier of $Y$ given $X,A$. That is, for every $x\in\mathcal{X},a\in\mathcal{A}$, $f^{*}(x,a)$ is the conditional distribution of $Y$ given $x,a$, i.e. $f^{*}(x,a)=\mathbb{P}\left(Y=\cdot|X=x,A=a\right)$. Denote by $(X^{\prime},A)$ a new pair of random variables, taking values in $\Delta_{\mathcal{Y}}\times A$, given by $(X^{\prime},A)=(f^{*}(X,A),A)$.
Lemma 3.2 (Factorization).
For any representation $Z$ of $(X,A)$, there is a representation $Z^{\prime}$ of $(X^{\prime},A)$, such that
$\mathbb{E}_{z^{\prime}\sim Z^{\prime}}h(\mathbb{P}\left(Y|Z^{\prime}=z^{\prime}\right))\leq\mathbb{E}_{z\sim Z}h(\mathbb{P}\left(Y|Z=z\right))\text{ and }D_{TV}(Z^{\prime})\leq D_{TV}(Z).$ (9)
In words, for every representation $Z$, we can find a representation $Z^{\prime}$ that only accesses $(x,a)$ through the value $f^{*}(x,a)$, and is at least as good in terms of both fairness and performance. Equivalently, this means that any two points $(x_{1},a)$ and $(x_{2},a)$ with coinciding conditional $Y$ distribution may be treated as identical for the purposes of constructing optimal representations. As a result, to find optimal tradeoffs, we can only consider the representations $Z^{\prime}$ on the small space $\Delta_{\mathcal{Y}}\times\mathcal{A}$, rather than $Z$ on the much bigger space $\mathcal{X}\times\mathcal{A}$.
Observations related to Lemma 3.2 were made in the context of classification in Kim et al., (2020),Xian et al., (2023), and Wang et al., (2023), which also start from the Bayes optimal classifier. Lemma 3.2 generalizes these observations to representations and to general losses. The proof may be found in Supplementary K.
3.3 The Invertibility Theorem